OpenStudy (anonymous):

What is the vertex for the function? (You will want to find the axis of symmetry first.) Write as an ordered pair, leaving no spaces, such as (1,-4). y = -3x2 + 6x

OpenStudy (anonymous):

@Luigi0210 @mathstudent55 @thomaster

OpenStudy (anonymous):

@shkrina

OpenStudy (anonymous):

do u have any guess ..

OpenStudy (anonymous):

idk how to find the axis of symmetry so no

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4f57de86e4b0c9fd2d4b7d90 check this out .... its explained here

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

u did not get that its same sum ...

OpenStudy (anonymous):

anyways @amistre64

OpenStudy (anonymous):

it didnt make sense

OpenStudy (anonymous):

given: ax^2+bx+c the vertex is on the axis of symmetry: -b/2a when x= -b/2a we find y main points ,,,

OpenStudy (amistre64):

given the format: y = ax^2 + bx + c, the axis of symmetry that they vertex lives on is the value of -b/(2a) y = ax^2 + bx + c y = -3x^2 + 6x + 0 a = -3 b = 6 , so we have everything needed to define -b/(2a) the value of y for the vertex is simply achieved by plugging in the value of x that we determined

OpenStudy (amistre64):

OR, another method is to comlete the square to rewrite this into vertex form: y = m(x-a)^2 - b such that the vertext is (a,b)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so the answer is (-3,6) ?

OpenStudy (amistre64):

y = -3x^2 + 6x y = -3( x^2 -2x) y = -3( x^2 -2x +1 -1) y = -3( x^2 -2x +1) + 3 y = -3(x-1)^2 + 3

OpenStudy (anonymous):

okay so is the answer (-3,6)

OpenStudy (amistre64):

no

OpenStudy (anonymous):

-__- i suck at math

OpenStudy (amistre64):

y = ax^2 + bx + c y = -3x^2 + 6x + 0 a = -3 b = 6 , so we have everything needed to define -b/(2a) x = -6/(2*-3) = 1

OpenStudy (anonymous):

so what do i need to do to get the answer!!

OpenStudy (amistre64):

well, from what has been posted so far; we determined that x=1, solve for y when x=1 y = -3(1)^2 + 6(1) = -3 + 6

OpenStudy (anonymous):

so y = 3 ?

OpenStudy (anonymous):

so its (1,3)

OpenStudy (amistre64):

that does seem to be correct yes :)

OpenStudy (anonymous):

alright thanks!!

OpenStudy (anonymous):

The parabola for the function will point in which direction? y = x2 + 3x - 4

OpenStudy (anonymous):

last question

OpenStudy (amistre64):

"point" ?

OpenStudy (anonymous):

upward or downward

OpenStudy (amistre64):

since the x^2 term has a positive coefficient, it will have the same form as y = x^2

OpenStudy (anonymous):

so upward?

OpenStudy (amistre64):

y = x^2 looks like a "U", so it opens upwards. if I were to take an educated guess, id assume they are thinking that it "points" downward like a "V"

OpenStudy (anonymous):

oh alright thanks again!

OpenStudy (amistre64):

good luck